### v1

# On-Chip Frequency Noise Cancellation in Nanomechanical Resonators using Cavity Optomechanics

Abstract
Understanding and minimizing the sources of frequency noise in nanomechanical resonators is crucial for many sensing applications. In this work, we report an ultracoherent perimeter-mode nanomechanical resonator co-integrated with an on-chip optical cavity. This device combines low thermomechanical force noise and low detector noise, allowing us to study its intrinsic frequency fluctuations in detail. We find that the fluctuations of two mechanical modes are strongly correlated. Moreover, we demonstrate the generation of a signal at the frequency difference between the two modes directly on chip via nonlinear optomechanical transduction. This ‘difference signal’ has vastly reduced intrinsic frequency fluctuations and can be used for frequency tracking with high precision, as we establish in a proof-of-principle experiment.
preprint: APS/
I Introduction
Precise tracking of mechanical eigenfrequencies is central to most applications of micro- and nanomechanical systems, including clocks, filters, and accelerometers [1, 2], as well as sensors for scanning force microscopy [3, 4, 5, 6], mass spectroscopy [7, 8], pressure sensing [9], and photothermal detection of nanoparticles [10]. In such applications, the measurement precision is limited by several sources of uncertainty [11, 12, 13, 14]: First, the thermomechanical motion of a resonator creates phase noise [11, 15] in the oscillation with a power spectral density (PSD) equal to $S_{\phi}^{\mathrm{th}}(\Omega)$, where $\Omega$ is the sideband frequency, i.e., the offset from the resonance frequency $\omega_{0}$. When measuring $\omega_{0}$, this phase noise leads to a ‘flat’ frequency noise PSD $S_{\omega}^{\mathrm{th}}(\Omega)\propto\Omega^{0}$. This noise sets a fundamental limit for the precision of frequency estimation at all time scales. Second, uncertainty in the frequency estimation is added by the readout (detection) noise. This noise leads to an apparent frequency noise $S_{\omega}^{\mathrm{det}}(\Omega)\propto\Omega^{2}$, which typically dominates the estimation of $\omega_{0}$ at short timescales. Third, coupling of the resonator to two-level systems [16] induces intrinsic frequency jitter [15], while coupling to an environment with temperature fluctuations in turn causes frequency fluctuations [17] owing to a finite thermal capacitance [11]. Together, both sources lead to a frequency noise PSD of $S_{\omega}^{\mathrm{int}}(\Omega)\propto 1/\Omega^{\alpha}$, with $\alpha$ between $0.5$ and $2$. This real frequency noise typically dominates at long timescales.
Various strategies have been employed to optimize frequency stability and its estimation. The thermomechanical contribution $S_{\omega}^{\mathrm{th}}$ is minimized by designing devices with low dissipation, which was one of the central efforts in the nanomechanics community over the last decade [18, 19, 20, 21, 22, 23]. Reducing the readout noise $S_{\omega}^{\mathrm{det}}$ requires better detection sensitivity, which can be achieved with cavity optomechanics [24, 25, 26, 27, 28, 29]. By contrast, overcoming intrinsic frequency fluctuations $S_{\omega}^{\mathrm{int}}$ is more challenging, as their microscopic origin is less well understood and harder to manipulate. In addition, this contribution impacts the measurement on long timescales and is not amenable to time averaging. The influence of intrinsic frequency fluctuations can be reduced by subtracting the correlated fluctuations of various modes in post-processing [30]. However, such analysis is generally susceptible to uncorrelated detector noise and cannot always be applied to real-time measurements. Alternatively, one can drive a nonlinear resonator at an optical working point [31, 32, 33], which requires careful calibration and fine-tuning of the system. In general, it would be desirable to have a method for direct, on-chip cancellation of frequency fluctuations that requires no special calibration.
In this work, we test an optomechanical platform that addresses all three frequency noise sources simultaneously. The device is an ultralow-dissipation perimeter-mode nanomechanical resonator [22] co-integrated with an on-chip optical cavity [27]. With low thermomechanical and readout noise, we can study the intrinsic frequency fluctuations over several order of magnitude in $\Omega$. Importantly, we find that the fluctuations measured on two different mechanical modes are strongly correlated. Taking advantage of the system’s nonlinear optomechanical transduction, we generate a signal at the frequency difference of the modes directly on the chip. This signal has vastly reduced frequency fluctuations and can be used as a resource for precise frequency tracking. We demonstrate this functionality in a proof-of-principle frequency measurement experiment.
II Experimental platform
The optomechanical system is formed by co-integrating a polygon nanomechanical resonator [22] with a photonic crystal optical micro-cavity. The polygon resonator is suspended in the near-field of the cavity and its motion couples dispersively to the optical modes through the evanescent field. Figs. 1(a) and (b) show scanning electron microscope (SEM) micrographs of a device. The devices are fabricated out of $250\text{\,}\mathrm{nm}$ thick stoichiometric silicon nitride (Si3N4) with 1.1 $1\text{\,}\mathrm{GPa}$ deposition stress. The polygon resonator is partially thinned down to $100\text{\,}\mathrm{nm}$. The devices are then patterned using electron beam lithography and dry etching and finally released using a KOH wet process. The details of the design and fabrication of these devices are discussed in Ref. [27].
[…]
Figure 1(c) illustrates our optical detection scheme. Continuous-wave laser light from a Toptica CTL 1550 is employed to drive the photonic crystal cavity at a fixed cavity detuning. We probe a cavity resonance at $1552.6\text{\,}\mathrm{nm}$ with a total linewidth of $\kappa/2\pi$ = $7.66\text{\,}\mathrm{GHz}$, see Fig. 1(d). The cavity has a free spectral range of $4.25\text{\,}\mathrm{n}\mathrm{m}$, see SI. Its DC component is used to lock the laser frequency $5.97\text{\,}\mathrm{GHz}$ red-detuned from the cavity resonance, while the AC component encodes the nanobeam’s motion. To minimize gas damping, the chip is placed in a vacuum chamber at a pressure of about $10^{-7}$ mbar. A piezo actuator, clamped to the chip, is used to excite the mechanical modes of the polygon resonator.
When the laser is detuned on the side of the optical resonance, we can observe the mechanical motion through direct detection. Using a digital lock-in amplifier (Zurich Instruments MFLI), we measure the power spectral density (PSD) of the mechanical displacement driven by thermomechanical force noise, see Fig. 1(e). The peaks correspond to the thermal motion of the mechanical modes of the polygon resonator, as confirmed by FEM simulations (see insets). This PSD is calibrated in units of frequency fluctuations of the optical resonance, see SI for details. The mechanical modes shown in this frequency range correspond to the high-$Q$ fundamental out-of-plane (OOP) perimeter mode and a family of in-plane (IP) modes, also confined to the perimeter of the device. The two modes used in our experiments are the OOP mode with frequency $\omega_{1}$ = $1.070\text{\,}\mathrm{MHz}$ and the IP mode with frequency of $\omega_{2}$ = $1.092\text{\,}\mathrm{MHz}$. Using the ringdown technique, we measure quality factors of $Q_{1}=4\times 10^{7}$ and $Q_{2}=3\times 10^{5}$ for the OOP and IP modes, respectively. For these modes, we also calibrate the vacuum optomechanical coupling rates to be $g_{0,1}/2\pi=$5.2\text{\,}\mathrm{kHz}$\$ and $g_{0,2}/2\pi=$15.6\text{\,}\mathrm{kHz}$\$ (see SI for ringdown measurements and details of the calibration).
We employ a digital phase-locked loop (PLL) to measure the modes’ frequencies as a function of time. The frequency noise PSD of the OOP mode shown in Fig. 1(f) is measured using different input optical powers at the same detuning from the cavity resonance, shown as a dashed grey line in Fig. 1(d). Due to the high $Q$, the thermomechanical contribution is too small to be measured and the spectra can be fit well with a model composed of only the $1/\Omega$ and detection noise contributions. We clearly see power-dependent detection noise at large sideband frequencies, while $1/\Omega$-type fluctuations at low sideband frequencies are independent of power. We use the $A/\Omega$ component of the spectra to calculate the corresponding Allan deviation as $\sigma_{\omega}=\sqrt{2A\ln{(2)}}$, where $A$ is a constant (see SI for details). These values are shown in the inset of Fig. 1(f), showing the power independence. These measurements confirm that the frequency noise for long integration times (low sideband frequencies) is dominated by $1/\Omega$-type fluctuations that are not related to optical absorption.
Figure 2: Properties of signal at $\omega_{\mathrm{d}}=\omega_{2}-\omega_{1}$. Displacement PSD of demodulated around (a) $\omega_{\mathrm{d}}=\omega_{2}-\omega_{1}$ (b) $\omega_{1}$, (c) $\omega_{2}$, and (d) $\omega_{s}=\omega_{1}+\omega_{2}$. The mechanical modes at $\omega_{1}$ (red) and $\omega_{2}$ (blue) are driven with separate PLLs. (e)-(g) Spectograms of the difference signal and the displacements of the OOP and the IP modes. The color bar corresponds to the normalized PSD. (h) Double-sided frequency-PSD for $\omega_{1}$, $\omega_{2}$, and $\omega_{\mathrm{d}}$ spanning $150\text{\,}\mathrm{s}$ with a data rate of $14.3\text{\,}\mathrm{kHz}$. We use the PLLs for estimating $\omega_{1}$ and $\omega_{2}$, while $\omega_{\mathrm{d}}$ is calculated using phase-to-frequency conversion [37]. (i) Absolute Allan deviation $\sigma$ in  $\text{\,}\mathrm{Hz}$ calculated using the same traces as in (h).
III Nonlinear frequency noise cancellation
Interestingly, when $\omega_{1}$ and $\omega_{2}$ are driven simultaneously, we observe optical signatures at the sum and difference frequencies ($\omega_{s,d}=\omega_{2}\pm\omega_{1}$), see Fig. 2(a)-(d). These signatures are due to the non-linear optical transduction of the mechanical signal by the cavity (see supplementary for details) that leads to mixing of the two signals [38]. More precisely, the cavity detuning is a time-varying parameter due to the mechanical motion of the resonator. This modulates the intensity of the out-coupled cavity response, where the first-order signal is at the mechanical resonance frequencies and the second-order effects are at their linear combinations. The displacement PSDs in Figs. 2(a)-(d) provide direct access to various frequency signal components. However, the $1/\Omega$-type frequency fluctuations are not easily visible in such a PSD on short timescales. To visualize the frequency fluctuations, we monitor how the PSDs change over time, see Figs. 2(e)-(g). Indeed, we find that both $\omega_{1}$ and $\omega_{2}$ change by more than $30\text{\,}\mathrm{Hz}$—hundreds of linewidths for the OOP mode—over several hours. The fluctuations of the two modes are strongly correlated but exhibit no clear pattern as a function of time, as expected for $1/\Omega$-type fluctuations. We also observe strong dependence on temperature changes (see SI). By contrast, the signal at $\omega_{\mathrm{d}}$ is significantly more stable, with a maximum change of about $0.4\text{\,}\mathrm{Hz}$, roughly two orders of magnitude smaller than that of $\omega_{1,2}$. We illustrate this in the corresponding frequency noise PSD (with uniform distribution of points along x-axis [39, 40]) in Fig. 2(h). Here, the difference mode has much lower contributions at sideband frequencies below $1\text{\,}\mathrm{Hz}$, marked by a decrease by an order of magnitude in the $1/\Omega$-type branch. One also observes how the noise floor is limited by the thermomechanical noise contribution of the mode with lower $Q$, i.e. $\omega_{2}$ to around $1.4\text{\times}{10}^{-4}\text{\,}\mathrm{Hz}^{2}\mathrm{/}\mathrm{Hz}$. In the corresponding Allan deviation $\sigma$ in Fig. 2(i), $\sigma_{\omega_{1}}$ is at its minimum around $0.9\text{\,}\mathrm{s}$ at a value of $26.1\text{\,}\mathrm{mHz}$ and starts increasing again due to thermal drifts. The difference signal, however, has a $\sigma_{\omega_{\mathrm{d}}}$ as low as $7.5\text{\,}\mathrm{mHz}$ at $\tau=$25\text{\,}\mathrm{s}$\$. For integration times beyond roughly $1\text{\,}\mathrm{s}$, the frequency fluctuations are lowered by up to a factor of 20 in the difference signal. In all of these graphs, it is evident that the frequency difference exhibits lower fluctuations than the OOP and IP modes at longer integration times (low sideband frequencies). This effect is directly tied to the origin of this ‘difference signal’ as a mixing product of modes $\omega_{1}$ and $\omega_{2}$. As the frequency fluctuations of the two real modes are correlated, their frequency difference remains approximately constant.
IV Mass sensing simulated experiment
In Fig. 3, we demonstrate that the stable signal at $\omega_{\mathrm{d}}$ is potentially useful for frequency measurements. As in Fig. 2, we simultaneously drive and measure $\omega_{1}$ and $\omega_{2}$ with a PLL, while demodulating the signal at $\omega_{\mathrm{d}}$ without applied drive. Additionally, we use an active feedback loop tuned to induce a pure frequency shift $\Delta\widetilde{\omega}$ on the OOP mode at $\omega_{1}$ without affecting the mechanical linewidth (by applying a feedback in-phase with the oscillator position, resulting in softening the spring constant [41, 42, 43]). Toggling this feedback signal on and off results in frequency jumps of $\omega_{1}$ without affecting $\omega_{2}$. We refer to the frequency shift signal using the normalised gain parameter $U\in[0,1]$ where $U=1$ corresponds to the maximum $\Delta\widetilde{\omega}$ shown in the graph. This controllable frequency modulation is used to simulate the frequency jumps expected, for example, in single-particle mass sensing experiments [7, 8].
Figure 3: Measuring frequency jumps with high precision at $\omega_{\mathrm{d}}$. (a) Schematic illustration of the effect of frequency jumps $\Delta\omega$ acting on $\omega_{1}$, but not on $\omega_{2}$. (b) Expected frequency jumps $\Delta\widetilde{\omega}$ when a dispersive feedback on mode $\omega_{1}$ is toggled on and off. The two modes at $\omega_{1}$ and $\omega_{2}$ are driven by separate PLLs. Frequency shifts (c) $\Delta\omega_{1}$, (d) $\Delta\omega_{2}$, and (e) $\Delta\omega_{\mathrm{d}}$ are simulataneously measured as a function of time.
In the time trace of $\omega_{1}$ in Fig. 3, it is difficult to distinguish the response to $U$ by eye, as it is partially masked by frequency fluctuations. As expected, $\omega_{2}$ does not display any systematic response to $U$ at all. For a quantitative analysis, we refer to each section between the grey dashed vertical lines as a block. We calculate the signal-to-noise ratio (SNR) as our metric, where we use the mean frequency jump size between two blocks as the signal and the standard deviation within each block to quantify the noise. We compare the SNR of the measured $\omega_{\mathrm{d}}$ against the measured $\omega_{1}$ and the calculated $\omega^{\mathrm{cal}}_{\mathrm{d}}=\omega_{2}-\omega_{1}$. For unprocessed data, we obtain an SNR of $4.1\pm 0.3$ at $\omega_{\mathrm{d}}$ in contrast with an SNR of $2.4\pm 1.7$ at $\omega_{1}$ and an SNR of $1.3\pm 0.1$ for $\omega^{\mathrm{cal}}_{\mathrm{d}}$. If we perform a simple moving-average filter over 200 data points in time (corresponding to a sync filter time of $24\text{\,}\mathrm{ms}$), the SNR at $\omega_{\mathrm{d}}$ is $6.2\pm 1$, while we obtain an SNR of $2.5\pm 1.2$ at $\omega_{1}$ and $4.5\pm 0.6$ for $\omega^{\mathrm{cal}}_{\mathrm{d}}$. The SNR of the difference signal is consistently better than that of the post-processed $\omega^{\mathrm{cal}}_{\mathrm{d}}$, as well as that of $\omega_{1}$ itself. In the SI, we present an alternative analysis based on cross-correlations, arriving at a similar result.
We assign the improved SNR of the signal at $\omega_{\mathrm{d}}$ relative to that measured at $\omega_{1}$ to the cancellation of intrinsic frequency fluctuations at low $\Omega$. As we see in Fig. 2(h), this effect is significant up to roughly $\Omega=$1\text{\,}\mathrm{Hz}$\$, explaining the reduced frequency fluctuations and the improved SNR in Fig. 3. However, for $\Omega>$1\text{\,}\mathrm{Hz}$\$, we can see in Fig. 2(h) that the signal at $\omega_{\mathrm{d}}$ has a higher $S_{\omega}$ than $\omega_{1}$, due to the increased thermomechanical force noise that it inherits from $\omega_{2}$. This noise contribution can be mitigated by averaging the data over time. This is why the SNR at $\omega_{\mathrm{d}}$ increases from 4.1 to 6.2 when applying the moving-average filter, while the one at $\omega_{1}$ remains the same within the measurement uncertainty. In principle, the low-$\Omega$ fluctuations can also be reduced in post-processing by calculating the signal $\omega^{\mathrm{cal}}_{\mathrm{d}}=\omega_{2}-\omega_{1}$ [30]. However, the SNR of this signal always remains below that of $\omega_{\mathrm{d}}$ because it samples the detection noise twice, once when measuring $\omega_{1}$ and once for $\omega_{2}$. Assuming that the detection noise imprinted on the two PLL measurements is uncorrelated, the resulting high-$\Omega$ noise present for $\omega^{\mathrm{cal}}_{\mathrm{d}}$ should be roughly $\sqrt{2}$ times larger than that for $\omega_{\mathrm{d}}$. This is in agreement with the observation that averaging improves the SNR at $\omega^{\mathrm{cal}}_{\mathrm{d}}$ significantly. Nevertheless, it never quite catches up with the on-chip noise cancellation we demonstrate with $\omega_{\mathrm{d}}$.
V Conclusions and outlook
In this work, we perform frequency noise characterization of high-$Q$ devices and observe correlated intrinsic frequency fluctuations of different mechanical modes. This correlation for low mass [17], high-$Q$ resonators is expected to predominantly be correlated flicker [15], resulting from the shared defects interacting with the various mechanical modes. Using nonlinear optomechanical transduction, we show how the correlation can be used for on-chip frequency noise reduction. With this method, we achieve a 20 fold decrease in absolute Allan Deviation at timescales larger than 1 second. Furthermore, the method performs a factor $3.2$ better than off-chip frequency noise reduction by post-processing, attributed to the fact that the method is not limited by PLL errors in the frequency measurements of individual modes, as the detection noise is only sampled once, not independently for both modes. Finally, we verify the usability of this noise suppression method in a proof-of-principle experiment that mimics the conditions for mass sensing. We find similar behaviour across multiple devices and by engaging different modes. Therefore, we conclude that our method can be a useful resource for precise frequency sensing experiments, such as mass spectroscopy [7, 8] or photothermal particle detection [10]. High-precision frequency tracking, combined with the excellent force sensitivity and low detection noise of these perimeter-mode resonators, is also crucial for nuclear spin detection [44, 45], and ultimately for nanoscale magnetic resonance imaging [46].
In our demonstration, the limiting factor for frequency noise up to sidebands of $100\text{\,}\mathrm{Hz}$ is the thermomechanical force noise of the IP mode at $\omega_{1}$, whose quality factor is significantly lower than that of the OOP mode at $\omega_{2}$. Alternative resonator designs with two high-$Q$ modes [47, 22] can potentially overcome this problem and achieve even better performance. For instance, with two modes that both possess $Q=5\times 10^{7}$, the total measured frequency variance would be reduced from $134\text{\,}\mathrm{mHz}$ to $78\text{\,}\mathrm{mHz}$, with a frequency noise PSD below $3\text{\times}{10}^{-5}\text{\,}\mathrm{Hz}^{2}\mathrm{/}\mathrm{Hz}$ over sideband frequencies up to $100\text{\,}\mathrm{Hz}$. Such a device will exhibit consistently low frequency noise PSD ($S_{\omega}$) over a large range of sideband frequencies, serving as a precise time reference at both short and long timescales.
Acknowledgements.
This research has been possible thanks to the Swiss National Science Foundation (SNSF) through grant no. 200021_200412 and the Novo Nordisk Foundation through grant no. NNF22OC0077964. N.J.E. acknowledges funding from the European Research Council (ERC) under grant no. 101117144. We would like to thank Prof Albert Schliesser (NBI Copenhagen), Prof Silvan Schmid (TU Vienna), Hajrudin Bešić (TU Vienna) and Olivier Faist (Zurich Instruments) for insightful discussions and nudges; Diego Visani, Thomas Gisler, Nils Prumbaum, Urs Grob and Luis Mestre for in-lab assistance. The technical support for the vacuum-compatible fibre alignment stage from Walter Bachmann of the engineering office at Dept. of Physics, ETH Zürich is highly appreciated.
Appendix A Cavity Characterisation
We use a sharply tapered fibre, with a full tapering angle of about 6 $\deg$, for input coupling to the integrated optical cavity. We optimize the fibre position on top of the auxiliary waveguide and the polarization of the incoupled light before the optical circulator to maximize the reflected optical power. We then perform a wide-range wavelength scan of the light reflected back from the integrated optical cavity. We observe multiple cavity modes with a free-spectral range (FSR) of $4.25\text{\,}\mathrm{n}\mathrm{m}$. We select the mode at $1566\text{\,}\mathrm{nm}$ as the working mode. After using the fine scan function to record a high-resolution scan of the mode of interest, Fig. 4 (c) we lock to the red side of the cavity using the built-in cavity lock feature of the Toptica laser. The calibration of the x-axis is performed using a fibre-loop cavity (FLC) [48].
Figure 4: Wide cavity scan and narrow scan with calibration. (a) Wide-range reflection spectrum of the cavity showing multiple resonances with a FSR of $4.25\text{\,}\mathrm{n}\mathrm{m}$. (b) Reflected signal of a fibre loop cavity (FLC) used for calibrating the x-axis (c) Reflected power from the rightmost cavity mode in (a) as a function of laser detuning $\Delta/2\pi$. $\Delta=0$ indicates the cavity resonance. The fit in orange dashed yields a linewidth of $\kappa/2\pi=3.33$ GHz.
Appendix B Optomechanical Characterisation
We characterise the mechanical properties of the perimeter resonator by first measuring a thermal spectrum as shown in Fig. 5(a), at the red-detuned sideband of the optical cavity. The first and second order modes of the resonator as well as the fundamental perimeter mode alongside the FEM simulation of their mode shapes are shown in the figure. Fig. 5(b) shows an FEM simulation of the cross-section the optical mode in the waveguide in the presence of the nanobeam. From this simulation we can infer the derivative of the guided refractive index with respect to the gap (i.e. $dn_{\mathrm{eff}}/dX)$. One can calculate the optomechanical pulling factor using the relation $G=\omega_{c}\frac{1}{n_{\mathrm{eff}}}\frac{dn_{\mathrm{eff}}}{dX}$. For our device, we estimate this quantity to be about $0.5\text{\,}\mathrm{G}\mathrm{Hz}\mathrm{/}\mathrm{n}\mathrm{m}$ for the OOP perimeter mode, in agreement with our estimation of the optomechanical coupling rate.
Figure 5: Example of the frequency noise model fitting for extraction of the $1/\Omega$ component coefficient for the lowest value of the input power.
Following this, we drive the modes at $\omega_{1}$ and $\omega_{2}$ using a PLL feeding back on a piezo actuator. We then perform a ringdown for each mode to characterise the $Q$ of each mode as shown in main text Fig. 1(e).
Figure 6: Ringdown measurements of the perimeter resonator’s modes $\omega_{1}$ in red and $\omega_{2}$ in blue with their respective fits (dashed).
Appendix C Optical frequency noise calibration
In this section the calibration procedure for optical cavity frequency and vacuum optomechanical coupling rates is explained. Within the linear transduction regime, when the laser is detuned from the cavity resonance, in direct detection the photocurrent is given by
$I(t)=D^{(0)}+D^{(1)}\xi(t),$
where
$D^{(0)}$ $=1-\frac{4\eta(1-\eta)}{1+\bar{\delta}^{2}},$ (1)
$D^{(1)}$ $=4\eta(1-\eta)\frac{2\bar{\delta}}{(1+\bar{\delta}^{2})^{2}},$ (2)
and $\xi(t)$ is the cavity frequency noise normalized to $\kappa/2$, containing the mechanical displacement. See later in SI for the derivations. The technique used here is based on using the to use the DC component of the photocurrent to calibrate the AC component. The DC and AC components of the detected voltage is given by
$V_{\mathrm{DC}}$ $=Z_{\mathrm{DC}}D^{(0)},$ (3)
$V_{\mathrm{AC}}(t)$ $=Z_{\mathrm{AC}}D^{(1)}\xi(t),$ (4)
where $Z_{\mathrm{DC}}$ and $Z_{\mathrm{AC}}$ are values of the trans-impedance gain at DC and AC frequencies. For the power spectral density (PSD) of the AC signal, defined as $S_{VV}(\omega)=\int e^{i\omega\tau}\langle V_{\mathrm{AC}}(t)V_{\mathrm{AC}}(t+\tau)\rangle d\tau$, it can be written as
$S_{VV}(\omega)=\left[\frac{Z_{\mathrm{AC}}}{Z_{\mathrm{DC}}}\right]^{2}\left[\frac{V_{\mathrm{DC}}D^{(1)}}{D^{(0)}}\right]^{2}S_{\xi\xi}(\omega).$ (5)
Without knowing the exact value of the trans-impedance gains and only knowing the ratio $\frac{Z_{\mathrm{AC}}}{Z_{\mathrm{DC}}}$, one can fully calibrate the spectrum. Assuming the detector has a flat response and $\frac{Z_{\mathrm{AC}}}{Z_{\mathrm{DC}}}=1$ the conversion factor for voltage to detuning noise, $S_{\Delta\Delta}=C^{2}S_{VV}$, is given by
$C=\frac{\kappa}{2}\left[\frac{D^{(0)}}{V_{\mathrm{DC}}D^{(1)}}\right]\quad(\mathrm{Hz}/\mathrm{V})$
This conversion is used to calibrate the data presented in Fig. 1(e). To calibrate the vacuum optomechanical coupling rate $g_{0}$ one has to integrate both side of Eq. 5 around a mechanical sideband. Integration of $S_{\xi\xi}$ around a mechanical mode with frequency $\omega_{m}$ and OM coupling $g_{0}$ obtains $8g_{0}^{2}n_{\mathrm{th}}/k^{2}$ where $n_{\mathrm{th}}=k_{B}T/\hbar\omega$ is the thermal occupation of the mode. Using this, $g_{0}$ is given by
$g_{0}=\frac{CV_{\mathrm{rms}}}{\sqrt{2n_{\mathrm{th}}}},$ (6)
where $V_{\mathrm{rms}}$ is the RMS value of the mechanical sideband in the voltage spectrum.
Appendix D Frequency noise characterization
When measuring the mechanical frequency fluctuations , it is known that the measurement is contaminated by three main sources of noise: thermomechanical, detection, and intrinsic frequency noise. Together, they constitute the total added frequency noise, whose power spectral density is given by:
$S^{\text{tot}}_{\omega}(\Omega)=S_{\omega}^{\mathrm{th}}(\Omega)+S_{\omega}^{\mathrm{det}}(\Omega)+S_{\omega}^{\mathrm{int}}(\Omega).$ (7)
It is known that the thermomechanical term has no frequency dependence and that the detection term has a quadratic frequency scaling. For the data shown in the main-text Fig. 1(f), we use the model $S_{\omega}(\Omega)=B\Omega^{2}+A/\Omega$ for fitting and extracting the $1/\Omega$ noise component. We have not included a constant term since as mentioned in the main text, the thermomechanical component for the high-$Q$ OOP mode is too low to beresolved. In Fig. 7 we show an example of the full f-PSD, the segment used for fitting and the result of the fit for the lowest input optical power. The roll-off at frequencies above $1\text{\,}\mathrm{kHz}$ is a result of the finite bandwidth of the lock-in amplifier’s demodulator. We do not include this segment in the model. The initial points of the data are not included either due to the lack of a sufficient number of points and large variations. $1/\Omega$-type frequency noise with a frequency noise PSD $S_{\omega}(\Omega)=A/\Omega$ corresponds to the constant Allan variance with standard deviation $\sigma^{2}_{\omega}(\tau)=2A\ln{(2)}$. This relation is used to produce the inset of the main-text Fig. 1(f).
Figure 7: Example of the frequency noise model fitting for extraction of the $1/\Omega$ component coefficient for the lowest value of the input power.
We also observe the influence of driving power and optical power on the overall frequency noise. We observe that while both parameters influence the detector branch, the effect is differently observed. Tuning the drive power affects the overall signal strength while tuning the optical power reduces the overall noise floor, as we can correlate from the side peak around $1.2\text{\,}\mathrm{kHz}$ for the $\omega_{1}$ mode, which remains the same height above noise for changing drive power but becomes more pronounced with increasing optical power. On the second hand, decreasing optical power does not influence the thermomechanical noise floor but a decreasing drive power causes an increase in thermomechanical noise for the low-$Q$ mode at $\omega_{2}$.
Figure 8: Measured frequency power spectral density of modes $\omega_{1}$ and $\omega_{2}$ for different driving powers (left) and optical power at fixed drive power(right). Similar to the effect of varying the optical power circulating in the cavity, we see that the $1/\Omega$ noise branch remains consistent across all measurements, while the high-frequency response changes with the driving power.
Appendix E Optical Non-linear Transduction for difference Frequency generation
The presence of a difference (and sum) frequency signal in the reflected laser light comes from the non-linear transduction of the mechanical signal by the photonic cavity. To derive the said dependence, we begin by studying the photon-mechanics interaction. The equation of motion of the intra-cavity field $a(t)$ in an optical cavity of linewidth $\kappa$ at a detuning $\Delta$ is given by
$\dot{a}=\left(i\Delta(t)-\frac{\kappa}{2}\right)a+\sqrt{\kappa_{\mathrm{ext}}}a_{\mathrm{in}}$ (8)
We assume that the detuning is modulated by two mechanical modes driven to amplitudes $x_{1}(t)$ and $x_{2}(t)$ respectively
$\Delta(t)=\bar{\Delta}+G_{1}x_{1}(t)+G_{2}x_{2}(t)$ (9)
We further assume that the mechanical modes are strongly driven (compared to the thermomechanical drive), and hence
$x_{i}=X_{i}\cos\Omega_{i}t+X_{th}(t)\approx X_{i}\cos\Omega_{i}t$ (10)
and that we are well inside the sideband-unresolved regime, $\kappa\gg\Omega_{1},\Omega_{2}$. With this assumption, the optical field adiabatically follows the detuning fluctuations. The intra-cavity field can therefore be expanded as
$a(t)=\frac{2\sqrt{\kappa_{\mathrm{ext}}}a_{\mathrm{in}}}{\kappa}\frac{1}{1-i\frac{2\Delta(t)}{\kappa}}.$ (11)
We can define the normalized detuning, the cavity coupling efficiency and the normalised amplitude as
$\delta(t)$ $:=\frac{2\Delta(t)}{\kappa}$ (12)
$\bar{\delta}$ $:=\frac{2\bar{\Delta}}{\kappa}$ (13)
$\eta$ $:=\kappa_{\mathrm{ext}}/\kappa$ (14)
$\xi(t)$ $:=\frac{2G_{1}x_{1}}{\kappa}+\frac{2G_{2}x_{2}}{\kappa}$ (15)
Using the input-output relation
$a_{\mathrm{out}}=a_{\mathrm{in}}-\sqrt{\kappa_{\mathrm{ext}}}a,$ (16)
we can find the field reflected from the cavity
$a_{\mathrm{out}}=a_{\mathrm{in}}\left[1-\frac{2\eta}{1-i\delta(t)}\right]$ (17)
In a direct detection scheme, the photocurrent is expressed as
$I(t)=\frac{|a_{\mathrm{out}}|^{2}}{|a_{\mathrm{in}}|^{2}}=\left|1-\frac{2\eta}{1-i\delta(t)}\right|^{2}=1-\frac{4\eta(1-\eta)}{1+\delta^{2}}.$ (18)
and then expand $I(t)$ to second order in $\xi$
$\begin{split}I(t)\approx 1&-\frac{4\eta(1-\eta)}{1+\bar{\delta}^{2}}+\frac{4\eta(1-\eta)2\bar{\delta}}{(1+\bar{\delta}^{2})^{2}}\xi(t)\\ &+\frac{4\eta(1-\eta)(1-3\bar{\delta}^{2})}{(1+\bar{\delta}^{2})^{3}}\xi(t)^{2}\end{split}$ (19)
We define the transduction coefficients to different orders of $\xi$
$D^{(0)}$ $=1-\frac{4\eta(1-\eta)}{1+\bar{\delta}^{2}}$ (20)
$D^{(1)}$ $=\frac{4\eta(1-\eta)2\bar{\delta}}{(1+\bar{\delta}^{2})^{2}}$ (21)
$D^{(2)}$ $=\frac{4\eta(1-\eta)(1-3\bar{\delta}^{2})}{(1+\bar{\delta}^{2})^{3}}$ (22)
The photocurrent is given by
$\begin{split}I(t)=&D^{(0)}+\frac{D^{(2)}}{2}\left(\delta_{1}^{2}+\delta_{2}^{2}\right)+\\ &\frac{D^{(2)}}{2}\delta_{1}\delta_{2}\cos(\Omega_{2}-\Omega_{1})t+\\ &D^{(1)}\delta_{1}\cos\Omega_{1}t+D^{(1)}\delta_{2}\cos\Omega_{2}t+\\ &\frac{D^{(2)}}{2}\left\{\delta_{1}^{2}\cos 2\Omega_{1}t+\delta_{2}^{2}\cos 2\Omega_{2}t+\delta_{1}\delta_{2}\cos(\Omega_{2}+\Omega_{1})t\right\}\end{split}$ (23)
with $\delta_{i}=2G_{i}X_{i}/\kappa$. In particular, the difference frequency component is given by
$I_{\mathrm{diff}}(t)=\frac{D^{(2)}}{2}\delta_{1}\delta_{2}\cos(\Omega_{2}-\Omega_{1})t$ (24)
Since we are neglecting the thermomechanical noise (Equation 10), all the signal terms appear as delta functions in the spectrum and the power of the difference frequency sideband is given by
$\begin{split}P_{\mathrm{diff}}&=\frac{{D^{(2)}}^{2}\delta_{1}^{2}\delta_{2}^{2}}{8}\\ &=\frac{32\eta^{2}(1-\eta)^{2}G_{1}^{2}G_{2}^{2}X_{1}^{2}X_{2}^{2}}{\kappa^{2}}\left[\frac{1-3\bar{\delta}^{2}}{(1+\bar{\delta}^{2})^{3}}\right]^{2},\end{split}$ (25)
s shown in Fig. 9, one scans through the red-detuned sideband of the cavity and monitors a driven mode signal and the transduced signal difference. The dashed lines in Fig. 9(b) are a single free parameter fit, namely the signal amplitude. Evidently, the nonlinear transduction (the difference signal) vanishes at $\bar{\delta}=1/\sqrt{3}$, termed the magic detuning introduced in Ref:[38]. Including the thermomechanical noise would exactly replicate the Ref: [38].
Figure 9: (Above) Left-half (red-detuned side) of the optical cavity resonance shown in Fig. 4. (Below) Transduced signal for mode $\omega_{1}$ (High Q mode) and the signal at $\omega_{d}$ (Diff signal). the latter shows a decay in signal at the magic detuning, $\bar{\delta}=1/\sqrt{3}$ [38]
We also demonstrate the optical transduction nature of the signal with another experiment. The transduction difference signal is substantial when the mode frequencies $\omega_{1}$ and $\omega_{2}$ are both driven. In Fig. 10 we plot a symbolic signal ’Trigger’ which is unity only when both modes are simultaneously driven, and zero otherwise. While monitoring both mode amplitudes and the difference signal amplitude, the parameter of interest is the decay time of each signal. When the drive to $\omega_{1}$ is switched off, the decay rate of $\omega_{1}$ and $\omega_{\mathrm{d}}$ follow the same trend. However, if the same is done while driving $\omega_{1}$ and switching the drive for $\omega_{2}$ off, the decay rate of $\omega_{\mathrm{d}}$ is defined by that of $\omega_{2}$, hence demonstrating the pure transduction nature of the signal which exhibits no intrinsic decay rate of its own.
Figure 10: Simultaneous ringdown measurements. Both mechanical modes at frequencies $\omega_{1}$ and $\omega_{2}$ are driven simultaneously, and the corresponding response at the frequency difference $\omega_{\mathrm{d}}$ is monitored. In the first sequence, the drive at $\omega_{2}$ is turned off while $\omega_{1}$ remains driven. A correlated decay and subsequent revival of both the $\omega_{2}$ mode and the signal at $\omega_{\mathrm{d}}$ is observed, indicating coupling between these modes. In the second sequence, the drive at $\omega_{1}$ is turned off, and a similar correlated decay is observed between $\omega_{1}$ and $\omega_{\mathrm{d}}$.
Appendix F Data Acquisition and Processing
We employ two (Zurich Instruments) MFLI units for all mentioned measurements. The first lock-in is used for driving the mechanics of the two modes $\omega_{1}$ and $\omega_{2}$ using a demodulator and a PLL for each with target BW $2\text{\,}\mathrm{kHz}$ for each of the two modes, running at a data rate of $2.14\text{\times}{10}^{6}\text{\,}\mathrm{S}\mathrm{a}\mathrm{/}\mathrm{s}$. Two other demodulators are used for generating the feedback signal that softening and stiffens the spring constant of the mode $\omega_{1}$ used in the main-text Fig. 3. The proportional gain for the left and right sections is $20$ and $8$. The physics and implementation are almost identical to that of feedback cooling [41, 42], the only difference being that the feedback force $F_{fb}$ is not in phase with the velocity $dx/dt$ of the mechanical mode but with the position, $x$, of the mechanical mode, hence $\pi/2$ phase-shifted relative to that for feedback cooling. It therefore introduces a frequency shift $\Omega_{fb}$.
$m_{\mathrm{eff}}\frac{dx^{2}(t)}{dt^{2}}+m_{\mathrm{eff}}\Gamma_{m}\frac{dx(t)}{dt}+m_{\mathrm{eff}}\Omega_{m}^{2}x(t)=F_{\mathrm{dr}}(t)+F_{\mathrm{fb}}(t)$
$m_{\mathrm{eff}}\frac{dx^{2}(t)}{dt^{2}}+m_{\mathrm{eff}}\Gamma_{m}\frac{dx(t)}{dt}+m_{\mathrm{eff}}\left[\Omega_{m}^{2}\pm\Omega_{fb}^{2}\right]x(t)=F_{\mathrm{dr}}(t)$
Appendix G 2D drifts
Figure 11: Time evolution of the center frequency of the $\omega_{1}$, $\omega_{2}$ and $\omega_{\mathrm{d}}$ and the room temperature change over a period of 14 hours. The data shows a strong correlation in the temporal shifts of modes $\omega_{1}$ and $\omega_{2}$, while the signal at the frequency difference $\omega_{\mathrm{d}}$ remains stable and unaffected throughout the measurement period.
Measurement of these modes is done at room temperature. As seen in main text Fig. 2(e)-(g), the frequency changes suddenly around the $8\text{\,}\mathrm{h}\mathrm{r}$ mark. This is correlated with a change in temperature of the laboratory as shown in  Fig. 11 due to start of movement in the lab after a night of calm.
Appendix H Mass sensing demonstration: Non-linear intrinsic cancellation vs post-processing subtraction
We analyse the data shown in the main-text section IV. Mass Sensing Simulated Experiment using SNR as our figure of merit. Here, we present an alternative quantitative analysis. We refer to the correlations between any two signals $s_{1}(t)$ and $s_{2}(t)$ – we calculate their normalized correlation coefficient $\rho(s_{1},s_{2})=\langle s_{1},s_{2}\rangle/\sqrt{\langle s_{1}\rangle\langle s_{2}\rangle}$, where $\langle$$\rangle$ indicates an average over time, implying that 1 and 0 indicate fully correlated and uncorrelated signals, respectively. We refer to $U(t)$ as the feedback gain that causes an expected frequency shift $\Delta\widetilde{\omega}$. We use the non-averaged data for the below for immediately next discussion where the averaged data used for plotting the main-text Fig. 3 shows very similar numbers.
Calculating the normalised correlation signals, We find $\rho(U,\omega_{1})=0.53$ and $\rho(U,\omega_{2})=0.03$, confirming the visual impression. The correlated intrinsic frequency fluctuations of the two modes lead to a finite correlation $\rho(\omega_{1},\omega_{2})=0.44$ between the modes, independently of $U$.
[…]
Figure 12: Post-processed cancellation vs non-linear transduced signal for one-on-one comparison
Now we return to calculating this figure of merit as a function of averaging or smoothing of the raw data, by convolving the data with a box filter of varying lengths (https://stackoverflow.com/a/26337730). As expected, for all sampling rates, the transduced data outperforms the post-processed data. This also comes from the fact that $\omega_{2}$ is not a PLL-tracked signal but a phase to frequency calculated signal [37]. We perform the same experiment where we instead track $\omega_{1}$ and $\omega_{2}$ and calculate $\omega_{\mathrm{d}}$ from phase. Even in this case, in spite of being limited by the differential calculation which lowers $\rho$ at higher sampling rates, the transduced signal shows a better performance than the post-processed signal. In both cases, as the effective sampling rate approaches $10\text{\,}\mathrm{s}\mathrm{a}\mathrm{m}\mathrm{p}\mathrm{l}\mathrm{e}\mathrm{s}\mathrm{/}\mathrm{s}$ or lower, only the contributions from the $S_{\omega}$ branch in the f-PSD remain, hence their performance converges.
Figure 13: Normalised covariance between the transduced and post-processed $\omega_{d}$ as a function of averaging which translates into an effective sampling rate. The left-most datapoints corresponds to $837\text{\,}\mathrm{S}\mathrm{a}\mathrm{m}\mathrm{p}\mathrm{l}\mathrm{e}\mathrm{s}\mathrm{/}\mathrm{s}$, i.e. without any averaging.